OpenStudy (anonymous):

3. Solve the system by the substitution method. Show your work. 2y - x = 5 x2 + y2 - 25=0 ?

OpenStudy (anonymous):

y(4y + y - 20) = 0 so, y = 0 or 4y+y-20 = 0 5y -20 = 0 y =45y-20=05(4)-20=00=0 x = 2y -5when y=4x = 8-5 =32y - x = 5x2 + y2- 25 = 02y - x = 5 2y - 5 = x so x = 2y - 5 (2y - 5)² + y² - 25 = 0 (2y - 5)(2y - 5) 4y² - 10y - 10y + 25 4y² - 20y + 25 4y² - 20y + 25 + y² - 25 = 0 4y² + y² - 20y = 0 4y² + y² - 20y = 0 y(4y + y - 20) = 0 y=0 x= -5 2y - x = 5 2(0) - x = 5, so x= -5) (-5,0) this wwas my answer

OpenStudy (anonymous):

my teacher said 3. Solve for x using equation 1 and then substitute this in the second equation. Once you simplify you will have a quadratic equation which you can factor and solve. Remember to factor out the GCF first.. The solutions should be written as ordered pairs (x1,y1) and (x2,y2).

OpenStudy (anonymous):

@Abhisar @Clalgee @dan815 @ganeshie8

OpenStudy (anonymous):

if my answer is correct what did she mean she only gave me two out of five points for thisw question

OpenStudy (anonymous):

@ShadowLegendX

OpenStudy (anonymous):

@nevermind_justschool

OpenStudy (anonymous):

@peachpi

OpenStudy (anonymous):

For one, you should have two solutions

OpenStudy (anonymous):

or did I miss it?

OpenStudy (anonymous):

somone sadi to use (3,4)

OpenStudy (anonymous):

\(2y - x = 5\) → \(x=2y-5\) \(x^2 + y^2 - 25=0\) \((2y-5)^2+y^2-25=0\) \(4y^2-20y+25+y^2-25=0\) \(5y^2-20y=0\) \(5y(y-4)=0\) \(y=0,~y=4\) Back substitute to find x

OpenStudy (anonymous):

ok I see. On this line you needed to solve for both y's y(4y + y - 20) = 0 You got \(y = 0\), but missed \(4y+y-20=0\)

OpenStudy (anonymous):

wait there were two y's

OpenStudy (anonymous):

yep. That's a quadratic equation so it has two solutions

OpenStudy (anonymous):

ohh so i nned solve for the second one right

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

can you help me withthat im so sleepy but i needa get this done

OpenStudy (anonymous):

I already put it up there. It's 4

OpenStudy (anonymous):

ohh ik that thnx hahaha see my sleepiness is messing with my brain thnx again padre

OpenStudy (anonymous):

np

OpenStudy (anonymous):

how do i copy and paste that i cant seem to get it highlighted to copy it and its killing me lol

OpenStudy (anonymous):

help me one more time please @peachpi

OpenStudy (anonymous):

idk that happens a lot with chrome. try double clicking the whole thing then deleting what you don't want

OpenStudy (anonymous):

its mot working when i double click it it hightlights like 1 number

OpenStudy (anonymous):

maybe try another browser

OpenStudy (anonymous):

i got it highlighted but it wont let me coppy it

OpenStudy (anonymous):

is there another way you could write it for me or nah

OpenStudy (anonymous):

why can't you use what you had originally. It's right. Just add the line y = 4, x = 3 at the end

OpenStudy (anonymous):

ohh ok i thought i needed the whole thing sorry

OpenStudy (anonymous):

no, your stuff is good. you just missed one of the solutions, last step.

OpenStudy (anonymous):

ok thnx bro gn